612shelbi:

An event last 3 hours and 15 minutes. How long did the event last in seconds?

dude:

Okay to you first want to turn the hours to minutes 1 hour = 60 minutes 3 * 60 = 180 180 (3 hours to minutes)+15 (minutes) = 195 minutes total Now you want to convert the minutes to seconds 1 minute = 60 seconds 195*60 = 11,700 seconds

dude:

Or more simply \(\frac{3.15 hours}{}\cdot\frac{60 minutes}{\cancel{1 hour}}\cdot\frac{60 seconds}{\cancel{1 minute}}=>\)